# Where to Go Next Day: Multi-scale Spatial-Temporal Decoupled Model for Mid-term Human Mobility Prediction

### v1

@@ Abstract
Individual Human Mobility, Mid-term Prediction, Spatial-Temporal Model, Urban Computing, Epidemic Spreading Analysis
@@ I Introduction
Human mobility prediction is generally addressed at two scales: collective and individual [12]. This study concentrates on the individual scale, with a specific emphasis on predicting personal mobility patterns. The core challenge in individual mobility prediction revolves around next-location forecasting, which aims to predict an individual’s future destinations based on their historical movement data. Mobility trajectories are commonly represented as sequences of locations, and the primary task is to extract meaningful transition patterns from these sequences. By identifying regularities and trends in an individual’s movement, this approach seeks to enhance the precision of mobility predictions. Traditional models rely on Markov chains [13], while deep learning models apply advanced techniques such as Recurrent Neural Networks (RNNs) and Transformers [14, 15, 16]. Owning to the spatial characteristics of locations, researchers have explored specific modules to capture spatial relationships. In recent years, the advent of graph representation learning techniques has demonstrated promising potential, motivating the adoption of Graph Neural Networks (GNNs) to enhance the predictive performance [17, 18, 19, 20, 21].
@@
Figure 1: Task illustrations of next location prediction and mid-term prediction.
Next location prediction is widely applied for personalized services, such as recommending points of interest (POI), as illustrated in Figure 1(a). However, urban management applications such as travel management and epidemic transmission analysis necessitate predicting residents’ comprehensive trajectory for the next day. This form of prediction, referred to as mid-term prediction, emphasizes modeling complete daily travel behaviors, as depicted in Figure 1(b). To tackle the mid-term prediction task, typical next location prediction models often employ an iterative mechanism to forecast next-day travel behaviors. This involves using predicted outcomes alongside historical data to forecast the remaining trips on the following day. Despite these efforts, these methods present difficulties in achieving satisfactory results, as evidenced in our experiments. Consequently, we propose an effective solution tailored specifically for the mid-term prediction task.
@@
The characteristics of mid-term prediction are illustrated in Figure 2. Firstly, accurate prediction of an individual’s daily trajectory requires knowing their location at each time interval. However, a prolonged stay at a specific location results in consecutive repetitions of the same location within the location sequence. For instance, hourly observations often show repeated instances of home occupancy during the nighttime and early morning, as illustrated in Figure 2(a). Such consecutive location repetitions in trajectory sequences confound the sequence models in accurately identifying location transitions. Secondly, learning the daily travel patterns over multiple days is essential for accurately predicting future behaviors. Yet, the long location sequences pose challenges to the perception and memory capabilities of sequence models, as indicated in Figure 2(b). Thirdly, people’s daily travel activities exhibit multi-scale regular patterns such as daily recurrency and weekly periodicity. Daily recurrency describes the similarity of individual travel behavior day-to-day. Urban residents’ daily travel is typically concentrated in certain primary travel patterns. For instance, as depicted in Figure 2(c), approximately 60% of the individual’s workday trips involve commuting between home and work, while around 30% involve visiting other places after work before returning home. On the other hand, weekly periodicity signifies the cyclical travel patterns, for instance, similarities in travel behaviors between Mondays. These patterns are essential for the modeling of daily trajectories. Lastly, cities display diverse spatial relationships among their regions, encompassing proximity, population flows, and significant administrative connections, as exemplified in Figure 2(d). Multiscale spatial granularity provides diverse insights into urban structural features, aligning with the multiscale characteristics encompassing both short and long-distance travel. Investigating these multiscale spatial relationships can enhance location representations and improve model performance in downstream tasks.
@@
In this paper, we propose a Multi-Scale Spatial-Temporal Decoupled Predictor (MSTDP) for the mid-term prediction task. Our approach starts by decoupling individuals’ daily trips into separate location and duration chains, thereby eliminating consecutive repetitions and directly capturing the daily transition patterns. Following this, a two-layer hierarchical encoder is employed to extract multi-scale temporal patterns like daily recurrency and weekly periodicity, efficiently broadening the receptive field and capturing longer-term behavioral regularities. Subsequently, a Transformer-based decoder integrates encoded historical information and previously predicted partial results to forecast trajectories for the following day. Furthermore, we introduce a heterogeneous graph module to capture the spatial and accessibility relationships between regions at different scales, thereby enriching the semantic representation of locations and enhancing predictive accuracy. We evaluate MSTDP against nine baselines using five urban-scale mobile record datasets. Empirical results demonstrate the superior performance of the proposed model. We further conduct comprehensive statistical physics analyses using Boston as a case study, examining motifs, travel distances, origin-destination flows, and other dimensions to provide a multidimensional assessment of the model’s effectiveness. Moreover, we utilize the model’s predictions to analyze the epidemic transmission dynamics. Specifically, we simulate the epidemic spread using the SEIR model [22] and the population transfer probability matrices derived from travel trajectories. This enables monitoring of the cumulative new cases and the current active cases. Compared to the baseline, MSTDP demonstrated remarkable improvements with reductions in MAE by 87.6% for the cumulative new cases and 62.8% for the current active cases, highlighting the superior predictive capability of our model in modeling travel behaviors.
@@ II Related Work
Human mobility modeling typically employs sequential models to capture the transition relationships between locations. Early methods mainly rely on Markov Chain to construct location transition probability matrices, with techniques to incorporate temporal and spatial information to capture individual travel preferences [23, 13, 24]. The evolution of deep learning has effectively facilitated the modeling of complex interactions. The foundational framework of location sequence models has progressed from the RNN series [25, 14, 26] to attention-based Transformer architectures [16, 15], consequently enhancing the capacity to capture high-order transition patterns. In addition to the transition regularity, location sequences offer a wealth of spatiotemporal information. Thus, researchers have introduced innovative mechanisms from diverse perspectives such as temporal and spatial semantics to enhance sequence models [27, 28, 29, 30, 31, 32, 33, 34]. Besides, some work also incorporates location category information to further constrain the location selection [35, 36, 37]. However, these methods would encounter several challenges when applied to the mid-term mobility prediction task, as discussed in Introduction I. At present, there is limited research dedicated to mid-term behavioral prediction tasks, such as HTAED [38] and RLMob [39]. But both models still face difficulties including repeated consecutive locations and excessively long sequences, and they inadequately capture the multi-scale spatiotemporal patterns inherent in individual daily travel. To mitigate their limitations, we propose a spatiotemporal decoupled hierarchical framework, including a heterogenous graph learner, which enlarges the perception field and efficiently captures the multi-scale spatiotemporal periodic information.
Over the past few years, graph neural networks have enjoyed rapid development and have been increasingly applied to learn location representations [40, 18, 41, 42, 43]. Researchers customarily construct a flow graph and an adjacency graph, from which they derive two distinct location representations and devise mechanisms to effectively combine them [17, 44, 19]. In this work, we integrated these two graphs into a heterogeneous graph to learn an informative location representation. Additionally, our graph considers multi-level adjacency relationships and flow relationships to efficiently capture complex geographic spatial information.
@@ III-A Individual Trajectory
Analyzing individual mobility patterns throughout the day necessitates dense data. To this end, we leverage Call Detail Record (CDR) data, which records locations using latitude and longitude coordinates rather than Points of Interest (POI) as in LBSN data. To address privacy concerns and simplify the modeling process, researchers partition the urban area into grid cells of equal size and map individuals’ trajectories to cells. Each grid cell is treated as a discrete location visited by individuals [24]. The cell size used in this study is 1 km $\times$ 1 km. Here, the location set is denoted as =$\mathcal{L}=\{l_{1},\dots,l_{|\mathcal{L}|}\}$. We then partition each day into fixed-size timeslots. Using an hourly time window as an example, the daily trajectory is expressed as ${\rm D}=\{l^{0}_{p},\dots,l^{23}_{q}\}$, where $l^{t}_{p}$ denotes the stay location $l_{p}$ within the hour $[t,t+1),t\in\mathbb{N}$. Let $\mathcal{U}=\{u_{1},\dots,u_{|\mathcal{U}|}\}$ denotes a set of individuals. Then all of trajectories of an individual $u$ can be represented as ${\rm TRAJ}_{u}=\{{\rm D}_{1},\dots,{\rm D}_{\rm N}\}$, where ${\rm N}$ is the number of total trajectories of the user, ${\rm N}=|{\rm TRAJ}_{u}|$.
Task Formulation. Given an individual $u$ and his or her historical trajectories $\{{\rm D}_{1},\dots,{\rm D}_{k}\}$, the goal is to predict the next day trajectory $\hat{\rm D}_{k+1}=\{\hat{l}^{0}_{p},\dots,\hat{l}^{w}_{q}\}$, where $k\leq|{\rm TRAJ}_{u}|$1. $w$ is determined by the size of the time window. In the methodology section, we present the model’s details using hour-long windows as an example. In the experiments, different datasets require varying time window sizes, selected based on the level of data sparsity.
@@ III-B Basic Location Graph
Adjacency Graph is an undirected graph $G^{a}=(V,E^{a})$ to depict the proximity between locations. The node set $V$ aligns with the location set in the flow graph $G^{f}$. The undirected edge $e^{a}_{pq}$ in the edge set $E^{a}$ indicates that the geospatial distance between $v_{p}$ and $v_{q}$, denoted as $dist_{pq}$, is less than a pre-defined threshold $\bar{d}$. In our work, $\bar{d}$ is set to 2 km for a cell with 1 km width. $e^{a}_{pq}=1$ if $dist_{pq}\leq\bar{d}$ and $e^{a}_{pq}=0$ otherwise.
@@ IV-B Geospatial Embedder
Locations inherently possess spatial semantic information, inspiring the utilization of graph neural networks to learn location representation from the adjacency graph and flow graph [17, 44, 19]. However, these methods tend to learn embeddings independently from each graph and then merge the embeddings as the location representation, disregarding the interdependencies between the two graphs. Towards this, we introduce a heterogeneous graph, as illustrated in Figure 4, that combines both flow and adjacency relationships between locations. Furthermore, we expect to integrate multi-scale urban structural information into the graph. Considering the administrative divisions among regions, we build an admin-level graph with the administrative areas as nodes to enrich the spatial connections between locations. In this way, the heterogeneous graph considers both the cell-level and admin-level flow and adjacency relationships, along with the inclusion relationship between cells and administrative areas. The scale of administrative areas can be chosen as block, tract, town, etc.
@@
We formulate the heterogeneous graph as $G^{H}=(V_{\rm cell},V_{\rm adm},E^{f}_{\rm cell},E^{f}_{\rm adm},E^{a}_{\rm cell},E^{a}_{\rm adm},E^{i})$, where $V_{\rm cell}$ and $V_{\rm adm}$ are the node sets of cell and administrative area, $E^{f}_{*}$ and $E^{a}_{*}$ $(*\in[\rm cell,\rm adm])$ are the flow and adjacency edge sets, and $E^{i}$ is the inclusion edge sets. The undirected edge $e^{i}_{pq},e^{i}_{pq}\in E^{i}$, means that the cell $v^{\rm cell}_{p}$ is located in the administrative area $v^{\rm adm}_{q}$. The description of edges in $E^{f}_{*}$ and $E^{a}_{*}$ remain consistent with Preliminary III. Moreover, the flow edge in our study includes hourly population flow between connected nodes. The edge is defined as $\mathbf{e}^{f}_{pq}=\{f^{0}_{pq},\dots,f^{23}_{pq}\}$, where $f^{t}_{pq}$ means the number of trips from $v_{p}$ to $v_{q}$ during the time interval from the hour $t$ to $(t+1)$, $t\in[0,23)$.
The node embeddings incorporate neighbor information from various edges to enhance the semantic richness. We employ different message-passing mechanisms based on the characteristics of the heterogeneous edges. Due to the simplistic semantics of the adjacent edge and the inclusion edge, we apply GraphSage [46] as the backbone to model the message passing of these two types of edges. The graph layer is defined as:
@@
where $\mathbf{x}^{(n)}$ is node embedding in the $n$-th layer, $W^{1}$ and $W^{2}$ are learnable parameters, and “${\rm mean}$” is the mean operator to aggregate the neighborhood. The initial node features $\mathbf{x}^{(0)}$ comprises word embeddings of region IDs and the time-varying population count within the region. Regions refer to cells or administrative areas.
Considering the complexity of semantic information in the flow edges, we employ an edge-considered variant of Graph Attention Network (GAT) [47] for effective modeling:
@@
$\alpha_{pq}=\frac{exp(\sigma(\mathbf{a}^{T}[W_{p}\mathbf{x}_{p}||W_{q}\mathbf{x}_{q}||W_{pq}\mathbf{e}^{f}_{pq}]))}{\sum_{s\in N(p)\cup\{p\}}exp(\sigma(\mathbf{a}^{T}[W_{p}\mathbf{x}_{p}||W_{s}\mathbf{x}_{s}||W_{ps}\mathbf{e}^{f}_{ps}]))},$ (3)
@@
After $n$-layers heterogeneous graph propagation, we obtain the final cell representation as the learned location embedding. These learning processes can be summarized as:

### v2

@@ Abstract
Individual Human Mobility, Mid-term mobility prediction, Spatial-Temporal Model, Urban Computing, Epidemic Spreading Analysis
@@ I Introduction
Human mobility prediction is generally addressed at two scales: collective and individual [12]. This study concentrates on the individual scale, with a specific emphasis on predicting personal mobility patterns. The core challenge in individual mobility prediction revolves around next-location forecasting, which aims to predict an individual’s future destinations based on their historical movement data. Mobility trajectories are commonly represented as sequences of locations, and the primary task is to extract meaningful transition patterns from these sequences. By identifying regularities and trends in an individual’s movement, this approach seeks to enhance the precision of mobility predictions. Traditional models rely on Markov chains [13], while deep learning models apply advanced techniques such as Recurrent Neural Networks (RNNs) and Transformers [14, 15, 16]. Owing to the spatial characteristics of locations, researchers have explored specific modules to capture spatial relationships. Recent advancements in Graph Neural Networks (GNNs) [17, 18, 19, 20, 21, 22, 23] and Large Language Models (LLMs) [24, 25] have significantly enhanced human mobility prediction by capturing spatial-temporal dynamics and leveraging contextual sequential information. Their integration offers promising opportunities for more accurate and comprehensive modeling of mobility behaviors.
@@
Figure 1: Task illustrations of next location prediction and mid-term mobility prediction.
Predicting the next location is a widely studied task in human mobility analysis, primarily applied in personalized services such as point-of-interest (POI) recommendation and navigation assistance, as illustrated in Figure 1(a). These tasks focus on identifying short-term, localized movement patterns, such as determining an individual’s immediate next stop. Given their focus, next location prediction tasks typically rely on datasets with sparse and irregular temporal resolutions, which reflect the episodic nature of the movements being modeled. In contrast, urban management applications, such as travel demand forecasting and epidemic transmission modeling, require predicting comprehensive daily trajectories for residents. This task, referred to as mid-term mobility prediction, emphasizes modeling complete daily travel behaviors, as shown in Figure 1(b). Unlike next location prediction, mid-term mobility prediction is driven by the need to address broader societal challenges, such as optimizing urban planning, managing traffic flows, and assessing public health risks. To meet these objectives, it is essential to capture an individual’s full-day mobility patterns, which necessitates dense observation data capable of capturing sequential and structural continuity throughout the day. While typical next location prediction models have been adapted to mid-term mobility prediction tasks, they often rely on iterative mechanisms that predict one trip at a time by incrementally combining predicted outcomes with historical data to forecast subsequent trips. Existing approaches encounter significant obstacles in mid-term mobility prediction tasks, primarily due to their inability to effectively handle repeated consecutive locations and overly long sequences. Such issues not only challenge pattern recognition and memory capabilities but also result in compounding prediction errors over extended time horizons. Furthermore, these methods fail to incorporate multi-scale spatiotemporal patterns that are essential for capturing individuals’ daily travel behaviors in a holistic and realistic manner. Motivated by these limitations, we propose a novel framework specifically designed for the mid-term mobility prediction task, offering a robust and comprehensive solution to address the unique challenges posed by this application.
@@
The characteristics of mid-term mobility prediction are illustrated in Figure 2. Firstly, accurate prediction of an individual’s daily trajectory requires knowing their location at each time interval. However, a prolonged stay at a specific location results in consecutive repetitions of the same location within the location sequence. For instance, hourly observations often show repeated instances of home occupancy during the nighttime and early morning, as illustrated in Figure 2(a). Such consecutive location repetitions in trajectory sequences confound the sequence models in accurately identifying location transitions. Secondly, learning the daily travel patterns over multiple days is essential for accurately predicting future behaviors. Yet, the long location sequences pose challenges to the perception and memory capabilities of sequence models, as indicated in Figure 2(b). Thirdly, people’s daily travel activities exhibit multi-scale regular patterns, encompassing both daily recurrency and weekly periodicity. Daily recurrency refers to the consistency and similarity observed in individual travel behaviors across consecutive days. Urban residents’ daily travel often concentrates on a limited set of primary patterns, such as commuting between regular locations. For illustrative purposes, Figure 2(c) highlights a conceptual example, showing that a significant proportion of workday trips involve commuting directly between home and work, while a smaller proportion includes additional stops before returning home. It should be noted that the percentages (e.g., 60% and 30%) are not based on empirical data but are presented solely as illustrative examples to emphasize the regularity that underpins mobility patterns. On the other hand, weekly periodicity captures cyclical patterns in travel behaviors, such as the recurrence of similar journeys on Mondays. These regular patterns provide essential insights for the modeling of daily trajectories. Lastly, cities display diverse spatial relationships among their regions, encompassing proximity, population flows, and significant administrative connections, as exemplified in Figure 2(d). Multiscale spatial granularity provides diverse insights into urban structural features, aligning with the multiscale characteristics encompassing both short and long-distance travel. Investigating these multiscale spatial relationships can enhance location representations and improve model performance in downstream tasks.
@@
In this paper, we propose a Multi-Scale Spatial-Temporal Decoupled Predictor (MSTDP) for the mid-term mobility prediction task. Our approach starts by decoupling individuals’ daily trips into separate location and duration chains, thereby eliminating consecutive repetitions and directly capturing the daily transition patterns. Following this, a two-layer hierarchical encoder is employed to extract multi-scale temporal patterns like daily recurrency and weekly periodicity, efficiently broadening the receptive field and capturing longer-term behavioral regularities. Subsequently, a Transformer-based decoder integrates encoded historical information and previously predicted partial results to forecast trajectories for the following day. Furthermore, we introduce a heterogeneous graph module to capture the spatial and accessibility relationships between regions at different scales, thereby enriching the semantic representation of locations and enhancing predictive accuracy. We evaluate MSTDP against nine baselines using five urban-scale mobile record datasets. Empirical results demonstrate the superior performance of the proposed model. We further conduct comprehensive statistical physics analyses using Boston as a case study, examining motifs, travel distances, origin-destination flows, and other dimensions to provide a multidimensional assessment of the model’s effectiveness. Moreover, we utilize the model’s predictions to analyze the epidemic transmission dynamics. Specifically, we simulate the epidemic spread using the SEIR model [26] and the population transfer probability matrices derived from travel trajectories. This enables monitoring of the cumulative new cases and the current active cases. Compared to the baseline, MSTDP demonstrated remarkable improvements with reductions in MAE by 87.6% for the cumulative new cases and 62.8% for the current active cases, highlighting the superior predictive capability of our model in modeling travel behaviors.
@@ II-A Next location prediction
Existing work on individual mobility prediction primarily focuses on next location prediction, which typically utilizes sequential models to capture transition relationships between locations. Early methods mainly rely on Markov chains to construct location transition probability matrices, with techniques to incorporate temporal and spatial information to capture individual travel preferences [27, 13, 28]. The evolution of deep learning has effectively facilitated the modeling of complex interactions. The foundational framework of location sequence models has progressed from the RNN series [29, 14, 30] to attention-based Transformer architectures [16, 15], consequently enhancing the capacity to capture high-order transition patterns. In addition to the transition regularity, location sequences offer a wealth of spatiotemporal information. Thus, researchers have introduced innovative mechanisms from diverse perspectives, such as temporal and spatial semantics, to enhance sequence models [31, 32, 33, 34, 35, 36, 37, 38, 39]. Besides, some work also incorporates location category information to further constrain the location selection [40, 41, 42].
Recent advancements have also explored leveraging Large Language Models (LLMs) to capture rich contextual semantics [24, 25], adopting Test-Time Adaptation techniques to dynamically adjust the model based on test data [43], and utilizing hyperbolic space to model hierarchical relationships [44]. Moreover, some studies have incorporated crowd flow information to integrate group-level travel dynamics, further complementing individual mobility modeling [45, 46].
However, methods designed for next location prediction face several challenges when applied to the mid-term mobility prediction task. As discussed in Introduction I, the presence of repeated consecutive locations and excessively long sequences poses significant challenges to their pattern recognition and memory capabilities. Furthermore, these methods fail to effectively capture the multi-scale spatiotemporal patterns inherent in individuals’ daily travel behaviors. To mitigate their limitations, we propose a spatiotemporal decoupled hierarchical framework, including a heterogeneous graph learner, which enlarges the perception field and efficiently captures the multi-scale spatiotemporal information.
Over the past few years, graph neural networks have evolved rapidly and have been widely applied to study human mobility [47, 18, 48, 49, 50, 51]. Some approaches leverage user-POI heterogeneous graphs to model the interaction between individuals and locations, effectively capturing personalized mobility patterns and spatial dynamics [44, 23]. In contrast, other studies construct flow graphs and adjacency graphs to derive distinct representations of locations based on spatial-temporal relationships, devising mechanisms to combine them for enhanced prediction performance [17, 52, 19, 22]. In this work, we take a step further by introducing a heterogeneous graph that integrates flow, adjacency, and inclusion relationships to learn a more comprehensive and informative location representation. Unlike existing approaches, our method incorporates multi-level adjacency and flow relationships, enabling the efficient modeling of complex geographic spatial structures to better capture human mobility dynamics.
II-B Successive location prediction
Successive location prediction focuses on forecasting sequences of future visited locations based on historical trajectories combined with the known timing of location transitions [53, 54, 55]. Future transition times provide travel schedule information that simplifies the prediction task. However, such information is typically unavailable in real-world scenarios. In contrast, mid-term mobility prediction aims to infer future mobility patterns solely based on historical behaviors, without relying on predefined transition times. This task is application-driven, particularly in urban management contexts, and emphasizes time structures that align with habitual travel patterns, such as daily or weekly intervals. For example, predicting population travel behaviors for the next day is crucial for applications like traffic planning or epidemic control. Although some prior studies also focus on forecasting “next-day” mobility, they often broadly categorize these tasks under location prediction without explicitly defining or distinguishing mid-term mobility prediction as a unique problem [56]. In this work, we aim to provide a clear definition of mid-term mobility prediction, differentiate it from successive location prediction, and highlight its significance in supporting real-world applications.
@@ III-A Individual Trajectory
Analyzing individual mobility patterns throughout the day necessitates dense data. To this end, we leverage Call Detail Record (CDR) data, which records locations using latitude and longitude coordinates rather than Points of Interest (POI) as in LBSN data. To address privacy concerns and simplify the modeling process, researchers partition the urban area into grids of equal size and map individuals’ trajectories to grids. Each grid is treated as a discrete location visited by individuals [28]. The grid size used in this study is 1 km $\times$ 1 km. Here, the location set is denoted as $\mathcal{L}=\{l_{1},\dots,l_{|\mathcal{L}|}\}$. We then partition each day into fixed-size timeslots. Using an hourly time window as an example, the daily trajectory is expressed as ${\rm D}=\{l^{0}_{p},\dots,l^{23}_{q}\}$, where $l^{t}_{p}$ denotes the stay location $l_{p}$ within the hour $t,t+1),t\in\mathbb{N}$. Let $\mathcal{U}=\{u_{1},\dots,u_{|\mathcal{U}|}\}$ denotes a set of individuals. Then all of trajectories of an individual $u$ can be represented as ${\rm TRAJ}_{u}=\{{\rm D}_{1},\dots,{\rm D}_{\rm N}\}$, where $\rm N$} is the number of total trajectories of the user, ${\rm N}=|{\rm TRAJ}_{u}|$.
Task Formulation. Given an individual $u$ and his or her historical trajectories $\{{\rm D}_{1},\dots,{\rm D}_{k}\}$, the goal is to predict the next day trajectory $\hat{\rm D}_{k+1}=\{\hat{l}^{0}_{p},\dots,\hat{l}^{w}_{q}\}$, where $k\leq|{\rm TRAJ}_{u}|$ -1. $w$ is determined by the size of the time window. In the methodology section, we present the model’s details using hour-long windows as an example. In the experiments, different datasets require varying time window sizes, selected based on the level of data sparsity.
@@ III-B Basic Location Graph
Adjacency Graph is an undirected graph $G^{a}=(V,E^{a})$ to depict the proximity between locations. The node set $V$ aligns with the location set in the flow graph $G^{f}$. The undirected edge $e^{a}_{pq}$ in the edge set $E^{a}$ indicates that the geospatial distance between $v_{p}$ and $v_{q}$, denoted as $dist_{pq}$, is less than a pre-defined threshold $\bar{d}$. In our work, $\bar{d}$ is set to 2 km for a grid with 1 km width. $e^{a}_{pq}=1$ if $dist_{pq}\leq\bar{d}$ and $e^{a}_{pq}=0$ otherwise.
@@ IV-B Geospatial Embedder
Locations inherently possess spatial semantic information, inspiring the utilization of graph neural networks to learn location representation from the adjacency graph and flow graph [17, 52, 19, 22]. However, these methods tend to learn embeddings independently from each graph and then merge the embeddings as the location representation, disregarding the interdependencies between the two graphs. Towards this, we introduce a heterogeneous graph, as illustrated in Figure 4, that combines both flow and adjacency relationships between locations. Furthermore, we expect to integrate multi-scale urban structural information into the graph. Considering the administrative divisions among regions, we build an admin-level graph with the administrative areas as nodes to enrich the spatial connections between locations. In this way, the heterogeneous graph considers both the grid-level and admin-level flow and adjacency relationships, along with the inclusion relationship between grids and administrative areas. The scale of administrative areas can be chosen as block, tract, town, etc. In cases where administrative divisions are unavailable, coarser-grained grids can be used as a substitute for administrative regions.
@@
We formulate the heterogeneous graph as $G^{H}=(V_{\rm grid},V_{\rm adm},E^{f}_{\rm grid},E^{f}_{\rm adm},E^{a}_{\rm grid},E^{a}_{\rm adm},E^{i})$, where $V_{\rm grid}$ and $V_{\rm adm}$ are the node sets of grid and administrative area, $E^{f}_{*}$ and $E^{a}_{*}$ $(*\in[\rm grid,\rm adm])$ are the flow and adjacency edge sets, and $E^{i}$ is the inclusion edge sets. The undirected edge $e^{i}_{pq},e^{i}_{pq}\in E^{i}$, means that the grid $v^{\rm grid}_{p}$ is located in the administrative area $v^{\rm adm}_{q}$. The description of edges in $E^{f}_{*}$ and $E^{a}_{*}$ remain consistent with Preliminary III. Moreover, the flow edge in our study includes hourly population flow between connected nodes. The edge is defined as $\mathbf{e}^{f}_{pq}=\{f^{0}_{pq},\dots,f^{23}_{pq}\}$, where $f^{t}_{pq}$ means the number of trips from $v_{p}$ to $v_{q}$ during the time interval from the hour $t$ to $(t+1)$, $t\in[0,23)$.
The node embeddings incorporate neighbor information from various edges to enhance the semantic richness. We employ different message-passing mechanisms based on the characteristics of the heterogeneous edges. To model the adjacency and inclusion edges, which primarily encode straightforward topological relationships without additional attribute information, we employ GraphSAGE [58] as the backbone for message passing. GraphSAGE’s innovative neighbor-sampling and aggregation mechanisms are particularly well-suited for handling large-scale graphs where edge semantics are simplistic and the main focus is on capturing node adjacency relationships. Furthermore, its efficiency and scalability make it an effective choice for dynamically evolving network structures embedded within urban mobility systems. The graph layer is defined as:
@@
where $\mathbf{x}^{(n)}$ is node embedding in the $n$-th layer, $W^{1}$ and $W^{2}$ are learnable parameters, and $\rm mean$}” is the mean operator to aggregate the neighborhood. The initial node features $\mathbf{x}^{(0)}$ comprises word embeddings of region IDs and the time-varying population count within the region. Regions refer to grids or administrative areas.
For flow edges, which capture rich semantic information such as human mobility flows and their associated attributes, we employ an edge-considered variant of the Graph Attention Network (GAT) [59]. GAT leverages attention mechanisms to adaptively weight and aggregate edge features, enabling precise modeling of edge-level semantics in spatial-temporal contexts. This capacity for dynamic weighting makes GAT particularly effective for capturing nuanced relationships expressed through flow data in urban networks. This is formulated as:
@@
[…]
@@
After $n$-layers heterogeneous graph propagation, we obtain the final grid representation as the learned location embedding. These learning processes can be summarized as:
@@
It is important to note that while GraphSAGE and GAT are employed in this framework for their suitability to specific edge types, the proposed framework is inherently modular. The GNN components can be replaced with alternative architectures as advancements in graph neural network research continue to emerge. This flexibility ensures that the framework remains adaptable and scalable, allowing integration of state-of-the-art techniques to further enhance the modeling of spatial-temporal mobility dynamics in urban systems.